A manufacturer claims that the average weight of a tin of baked beans in 440g and the standard deviation is known to be 20g. From a random sample of 100 cans you calculate the average weight to be 435g. At the 95% level of confidence, is there a change in the average weight of a can of baked beans?


1
Expert's answer
2021-04-15T06:51:25-0400

"\mu = 440 \\

\sigma = 20 \\

n = 100 \\

\bar{x} = 435 \\

α = 0.05 \\

H_0 : \mu = 440 \\

H_1 : \mu ≠ 440"

Test concerning averages

"z = \frac{\bar{x}- \mu}{\sigma / \sqrt{n}} \\

z = \frac{435-440}{20/ \sqrt{100}} = -2.5"

Critical value at 0.05 significance level with 99 d.f. z = ±1.96

Critical regions: Two-tailed test. Reject H0 if z ≤ -1.96 or z≥1.96

Conclusion: Since z = -2.5 is less than -1.96, reject the null hypothesis at the 95% level of confidence. There is a change in the average weight of a can of baked beans.